"Webserver/e-campus migration and phone reboot"
We are continuing to upgrade systems this week and have more maintenance that we need to do. In this case we will be moving our webservices to new hardware this afternoon. The web server will appear to stay operational the entire time as far as viewing web pages will be concerned but publishing new content to the web server will be frozen at 1:00PM today and will remain locked until 6:00PM.
This will disable the following services during that window:
Central Supply Store
O&M Workorder system
In addition we will be rebooting the analog (dorms, faxes, elevators, etc) phones again this evening at 5:00PM. This outage will take at most 15 minutes. The new IP phones will not be effected at all.
If you have any difficulties please do call the helpdesk at 7107.
Thanks much for your patience and understanding...
JD for DL